OpenStudy (anonymous):

solve 7 = y - 2 3/7 help please show work

OpenStudy (anonymous):

\[7 = y -2(\frac{ 3 }{ 7 })\] \[7 = y -(\frac{ (2*7) +3 }{ 7 })\] \[7 = y -(\frac{ 17 }{ 7 })\] \[y = 7 -(\frac{ 17 }{ 7 })\] find y

OpenStudy (hockeychick23):

ok so basically to find y you need to get y all by itself, so the first thing you want to do is add 2 3/7 to both sides, that will leave you with: 7 + 2 3/7 = y y= 9 3/7

OpenStudy (anonymous):

thankyou so much !! @hockeychick23

OpenStudy (anonymous):

thank u!! @sangya21

OpenStudy (anonymous):

Anytime :)